# Towards a more realistic evaluation of machine learning models for bearing fault diagnosis

### v1

@@ 3.3 Related works
Several studies in the literature have highlighted the issue of data leakage in bearing fault diagnosis, including [13], [5], [6], and [12]. These works propose new data-splitting strategies that primarily avoid segmentation-level leakage, but still allow bearing-wise leakage to persist (with the exception of [13]).
Hendriks et al. [5] proposed a bearing-wise splitting strategy on the CWRU dataset. Their methodology closely resembles ours, with one critical difference: healthy signals were included in both training and testing sets without applying consistent separation criteria, resulting in data leakage from the healthy bearings.
@@
Abburi et al. [6] also proposed an alternative split strategy aimed at reducing bearing-level information leakage on the CWRU dataset. The authors employed traditional machine learning models in a multiclass classification setting with three fault types (inner race, outer race, ball) and the healthy condition. Their results showed that the bearing-wise split consistently led to worse performance across all metrics (accuracy, precision, recall, and macro F1-score) when compared to a random split that included bearing information leakage. However, multiclass formulations such as this are unable to fully prevent leakage from the healthy class in the CWRU—a limitation we address through our proposed multi-label framework. To address the issues on [5, 6], we introduce a revised split strategy for CWRU, detailed in Section 4.2.3.
Another notable contribution is the study by [13], which proposed a bearing-wise split under a multiclass classification setting, referred to as a “part-to-part” approach. Using the KAt and CMTH datasets, the authors demonstrated a high correlation between signals originating from the same bearing—violating the i.i.d. assumption—and showed the substantial drop in accuracy when shifting from a condition-wise to a bearing-wise split, revealing the effects of a bearing-level leakage.
@@
A comprehensive investigation into dataset biases and evaluation protocols was conducted by [12], who proposed evaluation methodologies for widely used datasets such as CWRU, PU, MFPT, IMS, and UOC. Their approach focused on mitigating segmentation-level leakage, advocating for condition-wise splits across all datasets under a multiclass framework, with F1-macro reported as an auxiliary metric. However, the results remained overoptimistic due to residual bearing-level leakage.
@@ 4 Methodology
In this section, we first introduce our general methodology, including a bearing-wise split example using a generic dataset, the evaluation metric, and the hyperparameter tuning and model evaluation protocol. We then apply and specialize this methodology to three public bearing diagnosis datasets: University of Ottawa (UORED-VAFCLS), Paderborn University (PU), and Case Western Reserve University (CWRU). Finally, we describe the features, deep learning architectures, and data augmentation techniques used in our experiments.
@@ 4.1 General Methodology
The problem of detection and diagnosis of bearing faults can be formulated in many different ways, which can profoundly impact what a model can predict and how it is evaluated. Based on our findings regarding data leakage, we propose a methodology founded on a bearing-wise data splitting with the adoption of a multi-label classification framework to accommodate the potential for co-occurring fault modes and mitigate the effect of class imbalance. In a broader context, our proposed multi-label approach to fault detection and diagnosis is that of constructing specific detectors for each possible type of fault. Then, conventional fault detection amounts to simply verifying if any of the specific detectors gives a positive output, while fault diagnosis amounts to retrieving which specific detectors give a positive output. Additionally, this approach facilitates the use of sophisticated binary classification metrics, such as the Area Under the ROC Curve (AUC-ROC), which provide a threshold-independent and more nuanced evaluation of diagnostic performance compared to standard accuracy scores. A single summary metric, the Macro AUROC, is obtained by averaging the AUROCs of all individual detectors.
@@ 4.1.1 Data Splitting
Central to our methodology is a strict, bearing-level data partitioning strategy designed to prevent data leakage and ensure a valid assessment of model generalization. To illustrate this principle, we define a generic dataset construct, specified in Figure 2. This dataset comprises $B=15$ unique physical bearings, each characterized by one of three health states: healthy, inner race fault, or outer race fault.
@@
4.1.2 Evaluation metric
Although multiclass accuracy is a common metric in the cited literature, it has significant limitations for bearing fault diagnosis. Firstly, it is sensitive to class imbalance, a prevalent issue in fault diagnosis datasets. While balanced accuracy can mitigate this by weighting classes, it introduces another problem: the fault frequencies in benchmark datasets rarely mirror real-world conditions, where healthy states dominate until a fault develops. Secondly, the metric’s formulation assumes mutually exclusive classes, which inherently prevents the diagnosis of co-occurring faults. Creating additional classes for every fault combination is often impractical, as these events are too rare in datasets for a model to learn effectively. Finally, multiclass accuracy lacks a tunable decision threshold, making it impossible to optimize for specific operational needs, such as minimizing the False Negative Rate to avoid missed detections.
A key advantage of the proposed multi-label framework is its compatibility with evaluation metrics that are robust to class imbalance. By treating the detection of each fault type as an independent binary classification problem, we can employ fine-grained, prevalence-independent metrics. The primary metric selected for this work is the Area Under the Receiver Operating Characteristic curve (AUROC). The AUROC provides a threshold-independent measure of a classifier’s ability to distinguish between positive and negative classes, making it particularly suitable for bearing fault datasets where healthy samples often vastly outnumber faulty ones.
To obtain a single, comprehensive measure of overall model performance across all $F$ fault types, we compute the macro-average of the individual AUROC scores. This aggregate metric is calculated as
$\text{Macro AUROC}=\frac{1}{F}\cdot\sum_{i=1}^{F}\text{AUROC}_{i}.$
@@ 4.1.3 Hyperparameter Tuning and Model evaluation
  • Performance Estimation (CV): The model, using the selected hyperparameters, is retrained and evaluated across 100 distinct, randomly generated train-test splits. Although some test data in this stage may have been seen during hyperparameter tuning, the large number of disjoint splits dilutes the influence of any specific instance. The final reported performance is the average metric (e.g., Macro AUROC) over these 100 runs, providing a more stable and representative estimate of the model’s generalization ability.
@@ 4.2.3 CWRU bearing fault dataset
Figure 6: Bearing configurations used in the CWRU dataset. Each cell represents a specific acquisition setup containing two bearings—one on the drive-end and one on the fan-end. Fault types are denoted as follows: I for inner race fault, O for outer race fault, B for ball fault, and H for healthy.
In the division proposed by [5], faulty bearings with the same fault size are grouped, resulting in three different subsets (7, 14, 21 mils) that later are used to train and evaluate models. For example, when the subset with a fault size of 7 mils is used for training, the remaining subsets of 14 and 21 mils are used as test sets. This division prevents the occurrence of the same faulty bearings in both the train and test datasets, effectively addressing the data leakage issue. However, this solution is not entirely foolproof, as the dataset contains a single healthy bearing configuration, which must be split into the three previously described data subsets. In the [5] approach, this division is based on load, with healthy bearings at loads of 1, 2, and 3 HP being assigned to the 7, 14, and 21 mils subsets, respectively. This amounts to a condition-wise split of the healthy data, resulting in data leakage. Note that any division of the healthy bearing configuration on the CWRU dataset necessarily results in data leakage.
@@
A more realistic data partitioning strategy involves creating subsets in which multiple loads, fault sizes, types, and locations appear randomly. In this context, we propose a split methodology that selects signals from a random fault size configuration for each fault location–type pair. To prevent data leakage, signals from the same bearing configuration (e.g., the same physical bearing under specific conditions) are restricted to either the training or testing set, but not both. The selected configurations form the test set, while all remaining signals are allocated to the training set.
@@
In addition to this hold-out strategy, a 3-fold cross-validation procedure can be applied following the same selection principles, with the intent of using it to optimize hyperparameters. The first fold is generated by randomly selecting a fault size configuration for each fault location–type pair, as done in the hold-out scenario. The remaining two folds are then created by applying additional hold-out splits on the residual configurations, excluding the first fold. An illustration of this 3-fold setup is shown in Figure 7.
@@
Figure 8: Example of the two healthy-bearing split scenarios in our proposed methodology.
@@ 4.2.4 Summary
CWRU 20333Excluding bearings with a 0.028” fault size. 1 18444We only consider instances with a faulty bearing accompanied by a healthy bearing at the opposite side (see Section 4.2.3). 18 12
Paderborn 17555These bearings correspond to the group with real faults, excluding those with combined faults. 1 17 11 7
@@ 4.3.1 Shallow Learning models with bearing fault based features
Time + Frequency
RMS, peak-to-peak, kurtosis, skewness, crest factor,
magnitude of all bearing fault frequencies (1x - 5x)
on envelope spectra
@@ 5.2 PU dataset
Figure 10: Impact of train-test split ratio on model performance on the PU dataset using envelope spectrum as the input representation.
@@ 6 Experiments with data leakage
In this section, we investigate the effects of data leakage across multiple datasets by designing experiments that fit their unique structures. Earlier, we showed using a toy problem that data leakage produces overoptimistic results, which are not suitable for decision-making. We now present similar findings using real data.
@@ 7 Discussion and conclusion
Our experimental results demonstrate that when using a leakage-free methodology, deep learning models do not exclusively yield the best performance. In fact, shallow models trained with handcrafted features can be highly competitive and, in some scenarios, even superior. This is clearly illustrated by our findings on the CWRU dataset, where a Random Forest model achieved a Macro AUROC of 85.52% ± 7.43%, outperforming all deep learning baselines. Even when not the top performer, such traditional techniques provide a crucial performance baseline. Conversely, on the PU dataset, the WDCNN with spectrum envelope input yielded the best performance (80.16% ± 14.48%), highlighting the importance of choosing a good input representation. Finnaly, on the UORED-VAFCLS dataset, WDCNN with frequency-domain input outperformed all other configurations, reaching a Macro AUROC of 93.12% ± 4.26%.
@@
Based on these insights, we recommend the following to researchers in the field:
@@
  • Multi-label formulation: This formulation provides a setup for diagnosing co-occurring faults, a realistic scenario, and enables a more precise evaluation with prevalence-independent metrics, such as the Macro AUROC metric.

### v2

@@ 3.4 Related works
Several studies in the literature have highlighted the issue of data leakage in bearing fault diagnosis, including [5], [6], [7], [8] and [9]. Three of these works propose new data-splitting strategies that primarily avoid segmentation-level leakage, but still allow bearing-wise leakage to persist, with the exception of [6] and [8].
Hendriks et al. [7] proposed a fault-size splitting strategy for the CWRU dataset, considering it a better approach to obtain domain shift in bearing fault diagnosis. Since in the CWRU dataset each combination of fault size and type (i.e. inner, outer, ball) correspond to a unique bearing, splitting by fault size indirectly produces as a bearing-wise partition. However, signals from the single healthy configuration (consisting of a healthy bearing in the fan end and another healthy bearing in the drive end) were included in both training and test sets, resulting in data leakage.
@@
Abburi et al. [9] also proposed an alternative split strategy, now directly aimed at reducing bearing-level information leakage on the CWRU dataset. The authors employed traditional machine learning models in a multiclass classification setting with three fault types (inner race, outer race, ball) and the healthy condition. Their results showed that the bearing-wise split consistently led to worse performance across all metrics (accuracy, precision, recall, and macro F1-score) when compared to a random split that included bearing information leakage. However, they have also split the signals from the healthy configuration, leading to inflated model performance, notably on the binary fault detection (fault versus no fault) problem.
Matania et al. [8] highlighted common types of data leakage that can occur when using bearing fault diagnosis datasets, proposing a fault-size guided splitting strategy similar to [7], which they applied to the CWRU and PU datasets. Since this work primarily focused on diagnosis after a fault had been detected, it excluded healthy samples, thereby eliminating data leakage when utilizing CWRU. The paper claims that the correct way to avoid data leakage is to separate the data based on fault sizes. While this approach is suitable for CWRU and PU, where each bearing has a single fault size, it would not be applicable to datasets where each bearing is considered under multiple fault sizes (corresponding to the evolution of a fault over time), such as the UORED-VAFCLS. This limitation is clearly demonstrated in Section 6.1, where performance gains were observed in the bearing-level leakage experiment.
@@
A comprehensive investigation into dataset biases and evaluation protocols was conducted by [5], who proposed evaluation methodologies for widely used datasets such as CWRU, PU, MFPT, IMS, and UOC. Their approach focused on mitigating segmentation-level leakage, advocating for condition-wise splits across all datasets under a multiclass framework, with F1-macro reported as an auxiliary metric. However, the results remained overoptimistic due to residual bearing-level leakage, which was not addressed in their papers.
Another notable contribution is the study by [6], which proposed a bearing-wise split under a multiclass classification setting, referred to as a “part-to-part” approach. Using the KAt and CMTH datasets, the authors demonstrated a high correlation between signals originating from the same bearing—violating the i.i.d. assumption—and showed the substantial drop in accuracy when shifting from a condition-wise to a bearing-wise split, revealing the effects of a bearing-level leakage. Although this paper propose a similar analysis to ours, their experiments contain confounding factors that make it difficult to understand the origin of the decrease in model performance. In their work, data splitting strategies such as “run-to-run”, “day-to-day” and “part-to-part” were proposed. Each strategy correspond to completely different training set compositions, making it difficult to identify which one is more diverse. One could argue that the decrease in performance when changing to the “part-to-part” split is associated with the model being less diverse, rather than uniquely due to data leakage. Our paper focuses on solving this issue by proposing controlled data leakage experiments, where the training set remains fixed while varying the test sets.
@@ 4 Methodology
In this section, we first introduce our general methodology, which is exemplified using a hypothetical generic dataset; then, we apply and specialize this methodology to three public bearing diagnosis datasets: University of Ottawa (UORED-VAFCLS), Paderborn University (PU), and Case Western Reserve University (CWRU). Finally, we describe the features, deep learning architectures, and data augmentation techniques used in our experiments.
@@ 4.1 General Methodology
Our general methodology consists of three parts: the bearing-wise data splitting strategy; our problem formulation as multi-label binary classification with ROC-based evaluation metrics; and a hyperparameter tuning and model evaluation protocol chosen to minimize bias.
@@ 4.1.1 Data Splitting
Central to our methodology is a strict, bearing-wise data partitioning strategy designed to prevent data leakage and ensure a valid assessment of model generalization. To illustrate this principle, we define a generic dataset construct, specified in Figure 2. This dataset comprises $B=15$ unique physical bearings, each characterized by one of three health states: healthy, inner race fault, or outer race fault.
@@
4.1.2 Problem formulation and evaluation metric
The fundamental objective in bearing health monitoring is fault detection, the ability to reliably distinguish between normal and faulty operating conditions. While the secondary objective of diagnosis (identifying the specific fault mode) is critical, it relies heavily on the robustness of this initial detection. The predominant problem formulation in literature combines detection with diagnosis using a multiclass framework, in which the healthy condition is treated as one of the classes, alongside fault types. Under this formulation, accuracy is widely used as the evaluation metric, although it presents significant limitations in this context. Standard multiclass accuracy is an inadequate measure of detection quality due to the inherent class imbalance in typical fault diagnosis datasets, which contain a disproportionate number of faulty samples compared to healthy ones. Consequently, relying on a multiclass formulation can be misleading, as a model may achieve a high score despite classifying all healthy samples as one or more fault classes, failing to reflect performance in real-world scenarios where the healthy class is more prevalent.
One approach to mitigating the limitations of a multiclass framework involves decoupling the tasks of detection and diagnosis into a two-stage process. In the first stage, fault detection is treated as a standalone binary classification problem. To account for the characteristic class imbalance of these datasets, one might employ a prevalence-independent metric, as discussed in Section 2.2. In the second stage, a multiclass framework can be used for diagnosis, although it introduces significant practical drawbacks. The traditional multiclass formulation assumes that fault modes are mutually exclusive (only one can exist) and exhaustive (all possibilities are covered), which precludes the detection of co-occurring faults and complicates the handling of novel defect types. While it is possible to create classes that account for co-occurring faults, it is often unpractical and limited, specially due to the lack of samples in public datasets that correspond to these cases. Furthermore, accuracy is often a misleading metric because it lacks the nuance to reflect varying importance levels. In contexts where one fault class is significantly more damaging than others, accuracy fails to penalize critical misses more heavily than minor ones. This makes it difficult to optimize a model for scenarios where certain errors are far more costly than others.
To address these limitations and provide a simpler structure that combines detection and diagnosis, we advocate for a multi-label framework that treats each fault type (excluding the healthy state, e.g., Inner, Outer, Ball) as an independent binary classification problem. In other words, the model outputs a yes/no classification for each fault type, with the healthy state being understood as the case where no fault type is present. [^3] Naturally, this interpretation assumes that all possible fault types are included as classifier outputs. If additional unmodeled faults may occur, then the case of no fault detected should not be interpreted as a healthy state, but simply as absence of known faults. Under this framework, conventional fault detection amounts to simply verifying if any of the specific detectors gives a positive output, while fault diagnosis amounts to retrieving which specific detectors give a positive output. This approach enables a nuanced evaluation through the ROC curve, providing the ability to independently select a decision threshold for each classifier. In particular, operating points can be adjusted to meet application-specific requirements, such as prioritizing a high TPR for more critical faults. It also provides transparent, fault-specific insight into classifier behavior, making the reliability of each detector explicit. An additional benefit is that the multi-label approach allows for the use of faulty signals of one label (such as Inner) as true negatives for all other fault types (e.g. Outer and Ball), which is an advantage over the multiclass formulation, specially in cases where healthy signals may be scarce, such as the CWRU dataset.
While the ROC curve provides a detailed characterization of classifier behavior across all possible decision thresholds and is highly recommended for final evaluation, its interpretation can be impractical during model development, where concise indicators are typically preferred. The AUROC summarizes the information conveyed by a ROC curve into a single scalar value, as described in Section 2.2. In a multi-label framework, where each fault type produces its own ROC curve, we therefore adopt the Macro AUROC as the primary metric for model development, providing an overall view of performance across all classes for hyperparameter tuning and model comparison.
@@ 4.1.3 Hyperparameter Tuning and Model evaluation
  • Performance Estimation (CV): The model, using the selected hyperparameters, is retrained and evaluated across 100 distinct, randomly generated train-test splits. Although some test data in this stage may have been seen during hyperparameter tuning, the large number of disjoint splits dilutes the influence of any specific instance. The final reported performance is the Macro AUROC over these 100 runs, providing a more stable and representative estimate of the model’s generalization ability.
@@ 4.2.3 CWRU bearing fault dataset
Figure 6: Bearing configurations used in the CWRU dataset. Each cell represents a specific acquisition setup containing two bearings: one located at the drive-end (D) and the other at the fan-end (F). Fault types are denoted as follows: I for inner race fault, O for outer race fault, B for ball fault, and H for healthy.
In the division proposed by [7], faulty bearings with the same fault size are grouped, resulting in three different subsets (7, 14, 21 mils) that later are used to train and evaluate models. For example, when the subset with a fault size of 7 mils is used for training, the remaining subsets of 14 and 21 mils are used as test sets. Since, in the CWRU dataset, each fault size at each location corresponds to a unique bearing, this division effectively prevents the occurrence of the same faulty bearings in both the train and test datasets, incidentally addressing the data leakage issue. However, this solution is not entirely foolproof, as the dataset contains a single healthy bearing configuration, which is split into the three previously described data subsets. In the [7] approach, this division is based on load, with healthy bearings at loads of 1, 2, and 3 HP being assigned to the 7, 14, and 21 mils subsets, respectively. This amounts to a condition-wise split of the healthy data, resulting in data leakage. Note that any division of the healthy bearing configuration on the CWRU dataset necessarily results in data leakage.
@@
A more realistic data partitioning strategy involves creating subsets in which multiple loads, fault sizes, types, and locations appear randomly. In this context, we propose a split methodology that selects signals from a random fault size configuration for each fault location–type pair, which prevents data leakage as each pair corresponds to an unique bearing. The selected configurations form the test set, while all remaining signals are allocated to the training set.
@@
In addition to this hold-out strategy, a 3-fold cross-validation procedure is applied following the same selection principles, with the intent of using it to optimize hyperparameters. The first fold is generated by randomly selecting a fault size configuration for each fault location–type pair, as done in the hold-out scenario. The remaining two folds are then created by applying additional hold-out splits on the residual configurations, excluding the first fold. This procedure yields three train–test configurations, each using two folds for training and the remaining fold for testing in a 2:1 proportion. An illustration of this 3-fold setup is shown in Figure 7.
@@
Figure 8: Illustration of the two healthy-bearing split scenarios in the proposed methodology: training with healthy signals from the fan-end (left) and training with healthy signals from the drive-end (right). Green cells indicate samples used for training, whereas blue cells indicate samples reserved for testing.
@@ 4.2.4 Summary
CWRU 20444Excluding bearings with a 0.028” fault size. 1 18555We only consider instances with a faulty bearing accompanied by a healthy bearing at the opposite side (see Section 4.2.3). 18 12
Paderborn 17666These bearings correspond to the group with real faults, excluding those with combined faults. 1 17 11 7
@@ 4.3.1 Shallow Learning models with bearing fault based features
Time + Frequency RMS, peak-to-peak, kurtosis, skewness, crest factor, magnitude of all bearing fault frequencies (1x - 5x) on envelope spectra
@@ 5.1 UORED-VAFCLS dataset
To better understand the performance of each individual classifier in the frequency-domain WDCNN, we analyzed the ROC curves (Figure 9) by calculating the horizontal average [16] across 100 runs. For instance, at TPR = 90% (FNR = 10%), we achieve an average FPR = 6.53%, 7.59%, 9.08% and 46.30% for the diagnosis of inner, outer, ball and cage faults, respectively. As can be seen, most classifiers (Inner, Outer, and Ball) show strong performance, achieving AUROC values greater than 96.7%. In contrast, the Cage classifier showed a significantly decreased performance, yielding an AUC of 80.8% with a large variance of 13.7%, indicating a more challenging classification problem for this specific fault type.
Figure 9: ROC curves on the UORED-VAFCLS dataset for the frequency-domain WDCNN. The solid curves represent the horizontal average across 100 runs, while, for each curve, the shaded region represents one standard deviation.
It is worth noting that both the original ROC curves and their horizontal averages exhibit step-like behavior, with constant TPR values over ranges of FPR. Although the test set contains up to 240 segments, these are derived from only 24 underlying signals, thus many of these segments are highly similar and receive very similar scores, leading to a limited number of effective operating points. As a result, the lack of smoothness reflects the reduced diversity of the data.
@@ 5.2 PU dataset
Figure 11: Impact of train-test split ratio on model performance on the PU dataset using WDCNN with envelope spectrum as the input representation.
@@ 6 Experiments with data leakage
In this section, we investigate the effects of data leakage across multiple datasets by designing experiments that fit their unique structures. Earlier, in Section 3.2, we showed using a toy problem that data leakage produces overoptimistic results, which are not suitable for decision-making. We now present similar findings using real data.
@@ 7 Discussion and conclusion
Our experimental results demonstrate that when using a leakage-free methodology, deep learning models do not exclusively yield the best performance. In fact, shallow models trained with handcrafted features can be highly competitive and, in some scenarios, even superior. This is clearly illustrated by our findings on the CWRU dataset, where a Random Forest model achieved a Macro AUROC of 84.40% ± 9.31%, outperforming all deep learning baselines. Even when not the top performer, such traditional techniques provide a crucial performance baseline. Conversely, on the PU dataset, the WDCNN with spectrum envelope input yielded the best performance (80.16% ± 14.48%), highlighting the importance of choosing a good input representation. Finally, on the UORED-VAFCLS dataset, WDCNN with frequency-domain input outperformed all other configurations, reaching a Macro AUROC of 93.12% ± 4.26%.
@@
Based on these insights, we provide the following recommendations to researchers in the field:
@@
  • Multi-label formulation: This formulation provides a natural setup for diagnosing co-occurring faults and enables a more precise evaluation with prevalence-independent metrics, such as the Macro AUROC metric for model development/selection and the individual ROC curves for final evaluation and operating-point selection.